You have 2 solutions of liquid. One contains 10% water and the other contains 30% water. How many quarts of each should you mix in order to obtain 30 quarts of mixture that is 25% water?
Answer by checkley75(3666)

.3x+.1(30-x)=.25*30
.3x+3-.1x=7.5
.2x=7.5-3
.2x=4.5
x=4.5/.2
x=22.5 qts. of 30% water.
30-22.5=7.5 qts of 10% water.
Proof:
.3*22.5+.1*7.5=7.5
6.75+.75=7.5
7.5=7.5